OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (jdoe0001):

keep in mind that |x+1| really means +1(x+1) AND -1(x+1) so you'd end up with 2 values for "x", one from each of the equations

OpenStudy (jdoe0001):

\(\bf −5|x + 1| = 10 \implies -5\left[1(x+1)\right] = 10\\ \textit{and the equation of } -5\left[-1(x+1)\right] = 10\)